NARROW BAND, CW, CONTINUOUSLY TUNABLE, NEAR- TO MID-IR PPLN OPO FOR HIGH-RESOLUTION SPECTROSCOPY

\underline{A. SCHIFFMAN}, C. D. NABORS, B. ATHERTON, Coherent, Inc., 5100 Patrick Henry Drive, Santa Clara, CA 95054.

We are developing a high power, cw optical parametric oscillator for high resolution infrared spectroscopy. The OPO is based on periodically poled lithium niobate, pumped by a single frequency Nd:YAG laser, to produce single frequency light in the near- and mid-IR. The narrow band output is continuously tunable over its broad output range, enabling sub-Doppler spectroscopies of molecular gases to be obtained. Details of the OPO, and its prospective applications in IR spectroscopy, are discussed.